# Semantic Structure-Mapping in LLM and Human Analogical Reasoning

### v1

@@ Abstract
Analogical reasoning is considered core to human learning and cognition. Recent studies have compared the analogical reasoning abilities of human subjects and Large Language Models (LLMs) on abstract symbol manipulation tasks, such as letter string analogies. However, these studies largely neglect analogical reasoning over semantically meaningful symbols, such as natural language words. This ability to draw analogies that link language to non-linguistic domains, which we term semantic structure-mapping, is thought to play a crucial role in language acquisition and broader cognitive development. We test human subjects and LLMs on analogical reasoning tasks that require the transfer of semantic structure and content from one domain to another. Advanced LLMs match human performance across many task variations. However, humans and LLMs respond differently to certain task variations and semantic distractors. Overall, our data suggest that LLMs are approaching human-level performance on these important cognitive tasks, but are not yet entirely human like.
@@
language models , analogies , structure-mapping
@@ 1 Introduction
The recent advances of large language models (LLMs) have raised the question of whether LLMs can serve as useful cognitive models in the study of various aspects of human learning, cognition, and behavior [1, 2, 3]. One such recent debate has focused on whether LLMs acquire the ability to perform analogical reasoning as a by-product of their self-supervised learning objective [4, 5, 6, 7]. Analogical reasoning—the ability to align abstract structures between a source and target domain—is posited to play a central role in human learning and generalization, for example, our ability to reason efficiently in unfamiliar domains [8, 9]. Thus, the question of whether LLMs can reason analogically in a human-like way directly bears on their ability to serve as computational models of human behavior beyond just next-word prediction.
Recent work has focused on the ability of advanced LLMs to match human analogical reasoning performance on tasks that involve recognition of spatial and logical transformations in matrices [4] or detecting patterns in strings of letters or numbers [7]. For example, Mitchell [7] uses analogy tasks such as abcd:abce::ijkl:?? in order to test the extent to which LLMs and humans can recognize and generalize abstract structures and operations (in this example, ordered sequences and successor functions). Such studies have produced mixed results, with evidence suggesting that advanced LLMs achieve the same performance and even produce similar error patterns to those observed in humans [4, 6], but with doubts remaining about the robustness of LLMs’ abilities, particularly with respect to increasingly abstract and challenging domains [10].
Previous work has focused almost exclusively on analogies using abstract and arbitrary symbols, where structures are derived from symbols’ spatial positions in the text prompt, but the symbols themselves are unimportant. This leaves out questions about reasoning analogically over semantically meaningful symbols, such as words in natural language. This type of analogical reasoning, which we call semantic structure-mapping, requires mapping between semantic structure in one domain (e.g., the relationship between a dog and a puppy, or that a dog has four legs) and non-semantic (arbitrary) structure in the other domain (e.g., spatial position in the text prompt). This type of mapping is thought to play a crucial role in human cognition and development, such as in the language-analogical reasoning feedback loop proposed by Structure-Mapping Theory (SMT) [11]. Moreover, if LLMs are to provide insight into how humans perform certain cognitive functions, it will likely involve the role of distributional semantic learning [12, 13, 14] in the acquisition or representation of those functions. Therefore, we focus on investigating how humans and LLMs compare in tasks requiring semantic structure-mapping and assessing whether patterns differ from those observed on tasks involving only arbitrary symbols.
We design two experiments, focused respectively on the mapping of semantic structure (i.e., semantic relationships between symbols, such as relating the symbol dog to the symbol puppy) and semantic content (i.e., information attached to a symbol such as the knowledge that a dog has four legs). In each experiment, the subject (human or LLM) is presented with a set of left-hand terms (the source domain) and a corresponding set of right-hand terms (the target domain), with the final right-hand term omitted. The subject is asked to fill in this blank. An exact copy of our prompt and an example question is shown in Figure 1. We design multiple variants of such questions designed to probe structure-mapping that involves semantic structure and semantic content, respectively. We additionally design a series of control and distractor conditions—e.g., interleaving informative mappings (square => C C C) with uninformative ones (lime => X X X) in order to expose differences in the underlying mechanism.
Overall, the most advanced LLMs we tested match human performance across our primary conditions, even producing human-like error patterns. However, significant differences emerge in several control settings. Even the most advanced LLMs show more sensitivity than humans to information presentation order and struggle to ignore irrelevant semantic information that humans readily dismiss. Thus, our results contribute to the ongoing debate about analogical reasoning, corroborating both work arguing for impressive LLM performance [4, 15] and work highlighting important mechanistic differences between humans and LLMs [10, 16]. [^1] Code and data are available at
https://github.com/AnonymousReview123/Semantic_Structure_Mapping_Anon.
By presenting data on the unique role of semantic structure and content in analogical reasoning, we suggest differences remain in how LLMs and humans represent and map semantic structure, although this gap may be closing as models increase in size and incorporate more diverse training signals. We argue that this has important implications for studying cognitive development and the role of LLMs in this research going forward.
2 Methods
2.1 Experiment Details
2.1.1 Semantic Structure
Each subject was presented with a quiz, which is a sequence of four such questions generated using four sets of base domains and four sets of target domains selected such that a participant sees each base and target domain exactly once. Eight variants of the task were devised to investigate the influence of task variations as described above.
Questions are introduced with the prompt “We are conducting an experiment on general reasoning abilities. Below we will show you various words and drawings of each, after which you will need to complete the last drawing. Respond as concisely as possible with only the last drawing.” We use the term “drawings” to describe the elements in the target domain because it loosely encapsulates the idea of mapping between the source and target domains. In a similar way to how drawings serve as partial structurally isomorphic representations that depict a subject with varying degrees of abstraction [17], the elements in our target domains establish a space of relations that are isomorphic to those in the source domain. In some cases the term “drawing” is straightforwardly applicable, as when the capitalization of characters corresponds to the term for a mature animal. In other cases the use is strained, as when capitalization instead corresponds to a shape being symmetrical. The transparently liberal use of the term “drawing” is used to prime subjects to reason creatively while attending to the correspondence between source and target domains. The prompt’s lack of reference to analogical reasoning accesses pre-theoretic responses to the extent possible. For the same purpose the experiment is introduced to human subjects and LLMs as studying “general reasoning abilities.”
2.1.2 Semantic Content
Each condition (described in Table 4) contains two quizzes, with four questions per quiz. Unless otherwise stated, methodological details of the Semantic Content experiment match those of the Semantic Structure experiment.
The four conditions are divided into those that require numeric reasoning and those that do not. Within the numeric and non-numeric conditions respectively, one condition utilizes only one dimension of variation (referred to as “single-attribute”) whereas another adds a second dimension of variation (“multi-attribute”). This allows for comparing the relative performance of human subjects and models when the task is made to require compositional reasoning over layered transformations.
Questions were formatted like the following example:
@@ 2.1.2 Semantic Content
In this example, the number of symbols corresponds to a number-of-legs feature, and the usage of exclamation marks and asterisks corresponds to an egg-laying feature (or, alternatively, a mammal feature). The right-hand sequences of characters thereby encode properties of the entities denoted by the left-hand words. Given that humans are two-legged mammals, the correct answer here would be * *. In order to solve this task, the participant must understand both aspects of the information encoded in the right-hand terms and then construct the answer by generalizing to a new example.
2.2 Participants
2.2.1 LLMs
We run our experiments on the following LLMs: GPT-3 [18], GPT-4 [19], Pythia-12B [20], Claude 2 [21], Claude 3 Opus [22], and Falcon-40B [23]. All of the above are transformer-based LLMs trained primarily on a next word prediction objective.
GPT-3 consists of a 175B parameter model trained on text completion and finetuned to produce more coherent answers. The details of GPT-4 are not publicly known, but it is considered by some sources to be a mixture-of-experts (MoE) model consisting of numerous GPT-3-scale language models [24]. GPT-4, unlike GPT-3, supplements text-completion pretraining and finetuning with reinforcement learning from human feedback (RLHF) in order to better align model outputs with the expectations of a human user. The training of Claude 2 also includes RLHF, but its performance falls short of GPT-4. The more recent Claude 3 (in our case, the most advanced Opus version) is considered to approximately match GPT-4 performance in general. GPT-3 and -4 are developed by OpenAI, whereas Claude 2 and 3 are developed by Anthropic. Pythia-12B and Falcon-40B are open-weights LLMs trained on a text-completion objective and consist of 12B and 40B parameters respectively. Neither undergoes RLHF. Pythia-12B is developed by EleutherAI, and Falcon-40B is developed by the Technology Innovation Institute.
2.2.2 Human Subjects
We also test human participants on our experiments. Reported in the main text are results obtained from 194 (mostly undergraduate) University-Name University students (132 in the Semantic Structure experiment, and 62 in the Semantic Content experiment). The split of participants between experiments approximately matches the 9:4 ratio of experiment conditions. The number of participants by condition are as follows: Defaults 18, Distracted 18, Only RHS 18, Permuted Pairs 17, Permuted Questions 17, Random Finals 15, Random Permuted Pairs 6, Randoms 8, Relational 15, Categorial 16, Multi Attribute 16, Numeric 16, Numeric Multi Attribute 14. The Relational, Categorial, Multi Attribute, Numeric, and Numeric Multi Attribute conditions each have two quizzes while the remaining conditions each have four quizzes per condition. Subjects were assigned randomly to a single quiz from one condition without the re-use of subjects. Roughly the same number of participants were assigned to each condition, with the exception of the Random and Random Permuted Pairs conditions. These were together assigned roughly the expected number of subjects for a single condition due to their similarity.
2.3 Statistical testing
In each experiment, we are interested in the relative performance of human subjects and the best-performing models and how this depends on the particular experiment conditions. Differences between most models and human subjects are large and do not require statistical analysis, and so we focus our statistical analysis on the performance of GPT-4 relative to human subjects and Claude 3 relative to human subjects.
For each experiment and pair of subjects (human subjects and GPT-4, or human subjects and Claude 3) we fit a logistic model to the data with and without interactions between the subject type and the experiment condition. In all cases, the outcome variable is the un-aggregated per-question score achieved by a subject (either a 0 or 1), and the predictor variables are experiment condition (e.g. “Defaults” or “Permuted Pairs”) and subject type (e.g. “human subjects” or “GPT-4”). We use four likelihood ratio tests to assess whether the interaction between subject type and experiment condition is significant for a given pair of subjects within a particular experiment, as motivated by Glover [25]. In all four cases the interaction is significant, and so we use simple effects analysis to investigate the direction and significance of the effect of subject type within particular conditions.
For the semantic content experiment, we additionally perform a logistic simple effects analysis comparing the performance of a single subject type (human, GPT-4, or Claude 3) in compositional versus non-compositional conditions for the numeric and non-numeric cases respectively with the non-compositional condition as reference. For example, we assess the effect of the condition being Multi Attribute with Categorial as the reference condition for only the subject type Claude 3 (and likewise for the other two examined subject types).
Further details are provided in Sections A.1 and A.2 of the Appendix.
3 Results
3.1 Mapping Semantic Structure
We first design a set of experiments investigating the ability of LLMs and human subjects to map semantic structure in the source domain onto arbitrary, non-semantic structure in the target domain. In this set of experiments, our source domain (left-hand side) is a set of words which are assumed to possess some relational structure, and our target domain (right-hand side) is a set of strings related via non-linguistic string operations.
@@ 3.1 Mapping Semantic Structure
squarerectanglecircleovalC C Cc c cC Cc c
Figure 1: An example question (from the Defaults condition of the Semantic Structure experiment) with a representation of the structure-mapping solution below. The source domain is in blue and the target domain is in orange (for the provided elements) and yellow (for the inferred element).
3.1.1 Overall Performance
Human subjects perform well overall, obtaining accuracy between 0.4 and 0.9 across the various conditions. The most advanced LLMs that we test attain accuracies in the range 0.1-0.95 across conditions. This performance range is comparable to prior work on analogical reasoning over arbitrary symbols. For example, the results of human subjects on the “zero-generalization setting” studied by both Webb et al. [4] and Mitchell et al. [10] range from 0.2-0.8 in the former study and from 0.5-1.0 in the latter study. Similarly, results for LLMs (GPT-3, GPT-3.5, and GPT-4) across those conditions range from 0.1-1.0 in the two studies. Thus, our data suggest that analogies involving semantic structure-mapping are not inherently easier or harder than those which make use of arbitrary symbols.
Our Defaults condition consists of lexical items as a source domain and one of several string operation relations as a target domain. To investigate the robustness of performance metrics, we introduce three control conditions: (1) Permuted Questions, in which we present unaltered versions of the core task with varied question ordering; (2) Permuted Pairs, in which we alter the order in which the lines of the analogy are presented; and (3) Distracted, in which we interleave unrelated mappings between the lines of the target analogy. These conditions are shown in Table 1. We do not expect Permuted Questions to materially alter the task, but might see some effect of the Permuted Pairs and Distracted conditions, as they could make the relevant relations less transparent: see, for example, work on the blocking advantage in humans [26] and in LLMs [27].
Defaults Basic test of semantic structure-mapping square => C C C rectangle => c c c circle => C C oval =>

### v2

# LLMs as Models for Analogical Reasoning

@@ Abstract
Analogical reasoning—the capacity to identify and map structural relationships between different domains—is fundamental to human cognition and learning. Recent studies have shown that large language models (LLMs) can sometimes match humans in analogical reasoning tasks, opening the possibility that analogical reasoning might emerge from domain general processes. However, it is still debated whether these emergent capacities are largely superficial and limited to simple relations seen during training or whether they rather encompass the flexible representational and mapping capabilities which are the focus of leading cognitive models of analogy. In this study, we introduce novel analogical reasoning tasks that require participants to map between semantically contentful words and sequences of letters and other abstract characters. This task necessitates the ability to flexibly re-represent rich semantic information—an ability which is known to be central to human analogy but which is thus far not well-captured by existing cognitive theories and models. We assess the performance of both human participants and LLMs on tasks focusing on reasoning from semantic structure and semantic content, introducing variations that test the robustness of their analogical inferences. Advanced LLMs match human performance across several conditions, though humans and LLMs respond differently to certain task variations and semantic distractors. Our results thus provide new evidence that LLMs might offer a how-possibly explanation of human analogical reasoning in contexts that are not yet well modeled by existing theories, but that even today’s best models are unlikely to yield how-actually explanations.
@@
language models , analogical reasoning , cognitive science
@@ 1 Introduction
Analogical reasoning—the ability to identify and map structural relationships between a source and target domain—is posited to play a central role in human learning and generalization. It enables us to transfer knowledge from familiar domains to unfamiliar ones, recognize abstract patterns across superficially different contexts, and generate creative solutions to novel problems. Analogy is often credited with giving humans the ability to reason efficiently in unfamiliar domains [1, 2] and is considered by some to be among the primary predictors of intelligence and higher-order thinking [3, 4]. As a result, cognitive scientists have long sought to understand the computational mechanisms by which humans carry out this process.
With recent advances in artificial intelligence, there has been a growing interest in whether large language models (LLMs) possess a human-like capacity for analogical reasoning. Unlike existing computational models of analogy, LLMs contain no innate structure or inductive biases designed for analogical reasoning, and are not explicitly trained on analogical reasoning tasks. Rather, the behaviors that LLMs exhibit are the consequence of their relatively passive learning objectives, namely next-word prediction over large text corpora and, optionally, supervised fine-tuning for instruction-following and conversation. This raises the question of whether the capacity for analogical reasoning could emerge as a by-product of domain-general statistical learning, and if so, whether this could constrain theories of analogical reasoning in humans. Admittedly, LLMs do not represent a plausible account of how analogical reasoning might develop in humans. They require orders of magnitude more language data than children are estimated to receive before achieving fluency [5, 6] and typically rely on a developmentally unrealistic mix of text and computer code scraped from the internet. Even so, their success at analogical reasoning has important implications for cognitive theory: if analogical reasoning can emerge as a by-product of self-supervised learning and a generic neural architecture, it raises the burden of proof for computational and theoretical work on analogy which posits richer innate structure or more complex learning signals.
Work to date has yielded mixed interpretations of whether LLMs do in fact achieve human-level analogical reasoning. In 2023, Webb et al. found that GPT-3 could match or exceed typical human performance on most analogical tasks, including verbal analogies (e.g., ‘drive : car :: burn : ?’), letter-string analogies (e.g., transforming ‘a b c d’ to ‘a b c e’ and applying similar patterns to other sequences), and digit matrix problems (e.g., completing 3×3 matrices of digits following specific patterns or rules) [7]. Additionally, they found that GPT-4 showed strong performance on story analogies requiring the model to identify similar patterns between different stories, even when the stories described very different situations. Across the range of analogical reasoning problems tested, [7, 8] find that LLMs, in particular the most recent and advanced models, not only match human performance but replicate human-like error patterns as well.
However, contemporaneous work argues that LLMs’ abilities are overstated, and that their apparent analogical reasoning abilities might have alternative explanations. For example, LLMs’ performance degrades substantially when they are tested on stimuli that are deliberately designed to ensure that they are unlikely to appear in training data, whereas human performance remains stable [9]. It has yet to be determined whether this discrepancy is actually diagnostic of a failure of LLMs to engage in genuine analogical reasoning or whether it can be attributed instead to auxiliary task demands [10]. It thus remains an open question to what extent LLMs should be viewed as human-level or human-like with regard to analogy, and if so, whether LLMs’ success can advance cognitive theory in a meaningful way. The present state of this debate thus calls for novel analogical reasoning tasks that are likely not present in LLMs’ training data and which focus on aspects that have proven challenging for existing theoretical frameworks.
One such challenging aspect is flexible re-representation, the ability to dynamically restructure how concepts are encoded to support mapping between domains [11]. For example, a semantic content analogy from our stimuli is as follows:
@@
Completing this analogy requires inferring the abstract rule in which concepts on the left-hand side are represented in terms of two features: number-of-legs and is-mammal, with these features respectively determining the number and type of symbol which appears on the right-hand side. Since the concepts on the left-hand side have many features and it is not known at the outset which ones will be relevant to the mapping, a system for solving the task must possess the ability to dynamically adjust which features are used to represent the concepts. At the local level, individual concepts must often be re-represented in different ways depending on the analogy being drawn. For instance, the concept of ‘dog’ might need to be encoded in terms of its taxonomic category, physical attributes, diet, or domestication status. At a more global level, solving such tasks might invoke processes for schema induction [12] in order to recognize that superficially different scenarios instantiate the same abstract relational category or pattern.
While there exist a number of compelling cognitive theories of analogy, none readily account for the type of reasoning required by the above task, either because they do not address re-representation, or because they cannot be applied to real-world semantic concepts. For example, Structure Mapping Theory (SMT) emphasizes the importance of systematic relational mapping between structured representations [13]. While successful in explaining many aspects of human analogical reasoning, particularly the preference for mapping systems of relations rather than isolated features, SMT largely sidesteps the question of how appropriate representations are constructed. The theory has been criticized for assuming that concepts are already represented in predicate-logic form with clearly specified relations and thus failing to explain how humans identify which relations are relevant for a given analogy in the first place [11].
The first model which attempted to address the question of re-representation is Copycat [14]. Copycat was directly designed to account for the need to dynamically explore the space of possible representations in the course of analogy making and accomplished this by combining bottom-up and top-down processes. This facilitated re-representing the same letter string in multiple ways depending on the patterns being mapped. However, it has only been successfully applied to synthetic domains such as letter-string analogies. While the approach is intended to be domain-general in principle, in practice it would require significant theoretical and engineering effort to be adapted to the open domain concepts we employ here, if it is in fact possible to do so.
More recent approaches have attempted to bridge this gap. The BART (Bayesian Analogy with Relational Transformations) model accounts for the emergence of relational representations from non-relational inputs through statistical learning [15]. Other work has made progress on modeling re-representation in specific domains, such as visual patterns in Raven’s Progressive Matrices [16] or abstract concepts through program induction [17]. These models, like Copycat, include a set of bottom-up processes which can respond to top-down signals in order to account for re-representation in the course of analogy formation. Once again, these approaches remain limited to relatively constrained domains and have not been extended to handle open-ended semantic concepts.
This points to an important gap in our understanding of analogical inference: no current framework adequately explains how humans flexibly re-represent semantically rich concepts and relations to support novel analogical mappings and extract generalizable rules. There is thus a need for a computational model which can explain human behavior on complex analogical reasoning tasks using domain-general mechanisms.
To test whether LLMs might address this gap, we design a novel set of analogical reasoning tasks (split into two studies) which we present to both humans and LLMs. Our tasks focus respectively on reasoning from semantic structure and semantic content. Study 1 examines how subjects leverage semantic structure (relationships between concepts) to reason about arbitrary symbolic patterns in a target domain. Through systematic manipulation of available information—including conditions that remove linguistic content entirely or introduce misleading semantic relationships—we can isolate the role that semantic structure plays in guiding analogical inference. Study 2 extends this investigation by focusing on semantic content rather than just structural relationships. Subjects must identify relevant features of concepts (like number-of-legs and is-mammal) and map these onto abstract relations between arbitrary symbols. We manipulate the nature and number of attributes required to perform the analogical inference to assess how this affects performance in humans and LLMs.
Our experiments show that LLMs might provide a how-possibly model of non-trivial analogical reasoning in human subjects: the most successful LLMs achieve human-level performance on our challenging tasks which require abstract rule induction and re-representation. However, we also find that, while LLMs can reach human-level performance, even the best models demonstrate significantly different patterns of behavior in response to control conditions and stress tests, making them unlikely to yield how-actually explanations. In Study 1, we find that LLMs are more sensitive to presentation order and struggle when irrelevant semantic distractors are present. In Study 2, the best performing LLM matched human performance across all conditions, and exceeded it in the compositional variants on which human subjects performed most poorly.
Taken together, our findings suggest that LLMs can achieve sophisticated analogical reasoning capabilities through domain-general learning mechanisms, although they likely employ different mechanisms than humans. This suggests that LLMs are poised to offer insights on behaviors that current cognitive models cannot explain, and that research seeking to unpack the more precise mechanisms LLMs employ in order to achieve this behavior, while challenging, is likely to be a fruitful endeavor for cognitive theory.
2 Analogical Representation Tasks
2.1 Study 1: Analogical Inference from Semantic Structure
Study 1 investigates the capacity to map semantic relationships between concepts in a source domain to symbolic patterns in a non-semantic target domain. The task design [^1] The code and data used for generating all experiments are available at https://github.com/AnonymousReview123/Semantic_Structure_Mapping_Anon. specifically relates to the inference process in analogical reasoning, rather than to the initial mapping discovery that characterizes many traditional analogy tasks, and has a particular focus on how semantic structure can guide analogical inference in formal domains.
The core mechanism underlying the tasks in Study 1 involves a structural mapping where subjects must: (1) identify meaningful relations in the source domain, (2) abstract these relations into a form that can be mapped onto the target domain, and (3) generate appropriate completions by applying the mapped relations. Subjects are presented with several rows showing explicit mappings between words (source domain) and symbol sequences (target domain), followed by a final row containing only a word. Their task is to infer the corresponding symbol sequence that completes the pattern.
In the Defaults condition, subjects might see “dog” mapped to “X”, “puppy” mapped to “x”, and “cat” mapped to “Y”, and be asked to produce the mapping for “kitten”. To succeed at this task, subjects must first re-represent the concepts to identify relevant relational patterns (e.g., recognizing that both dog $\leftrightarrow$ puppy and cat $\leftrightarrow$ kitten instantiate an adult $\leftrightarrow$ juvenile relationship). Second, they must induce a general rule that maps this semantic relationship onto the formal pattern in the symbol sequences (e.g., adult $\rightarrow$ uppercase, juvenile $\rightarrow$ lowercase). A question from this condition is illustrated in Figure 1 along with the instructions shown to participants.
@@ 2.1 Study 1: Analogical Inference from Semantic Structure
Figure 1: An example question from the Defaults condition of the Semantic Structure experiment. A subject observes four questions in a row from one condition only with the answers to prior questions revealed (we call this sequence of four questions a “quiz”).
We then include several control conditions. The first set of these controls is intended to probe the reasoning mechanism, in particular testing how it responds to perturbations and distractions in how the questions are presented. Prior work suggests such changes could influence performance by making the relevant relations less transparent: see, for example, work on the blocking advantage in humans [18] and in LLMs [19].
The Permuted Pairs condition presents the same mappings in scrambled order, testing whether subjects can extract the underlying rule regardless of presentation sequence. This condition tests whether subjects’ performance is robust to the order in which example mappings are provided. While the Permuted Pairs condition changes the order of lines within each question to disrupt the analogy structure, the Permuted Questions condition leaves questions completely intact but shuffles the order in which the questions are presented in order to verify that the arbitrary question ordering does not have an unintended effect.
The Distracted condition introduces irrelevant mappings, requiring subjects to identify and suppress information that doesn’t fit the core analogical pattern. This condition tests whether subjects are able to ignore irrelevant mappings and rely exclusively on relevant ones for inference. The Defaults, Permuted Pairs, and Distracted conditions are shown in Table 1 below.
Defaults Basic test of analogical reasoning from a linguistic to a non-linguistic domain square => C C C rectangle => c c c circle => C C oval =>